### 05 Jan 2024 v2

# Knapsack: Connectedness, Path, and Shortest-Path

Abstract
We study the Knapsack problem with graph-theoretic constraints. That is, there exists a graph structure on the input set of items of Knapsack and the solution also needs to satisfy certain graph theoretic properties on top of the Knapsack constraints. In particular, we study Connected Knapsack where the solution must be a connected subset of items which has maximum value and satisfies the size constraint of the knapsack. We show that this problem is strongly $\mathsf{NP}$-complete even for graphs of maximum degree four and $\mathsf{NP}$-complete even for star graphs. On the other hand, we develop an algorithm running in time $\mathcal{O}\left(2^{\text{tw}\log\text{tw}}\cdot\text{poly}(n)\min\{s^{2},d^{2}\}\right)$ where $\text{tw},s,d,n$ are respectively treewidth of the graph, the size of the knapsack, the target value of the knapsack, and the number of items. We also exhibit a $(1-\varepsilon)$ factor approximation algorithm running in time $\mathcal{O}\left(2^{\text{tw}\log\text{tw}}\cdot\text{poly}(n,1/\varepsilon)\right)$ for every $\varepsilon>0$. We show similar results for Path Knapsack and Shortest Path Knapsack, where the solution must also induce a path and shortest path, respectively. Our results suggest that Connected Knapsack is computationally the hardest, followed by Path Knapsack and then Shortest Path Knapsack.
Keywords:
Knapsack Graph Algorithms Parameterised Complexity Approximation algorithm.
1 Introduction
The Knapsack problem is one of the most well-studied problems in computer science [5, 6, 18, 21]. Here, we are given a set of $n$ items with corresponding sizes $w_{1},\ldots,w_{n}$ and values $\alpha_{1},\ldots,\alpha_{n}$, the size (aka budget) $s$ of the bag, and the goal is to compute a subset of items whose total value is as high as possible and total size is at most $s$.
Often there exist natural graph structures on the set of items in the Knapsack problem, and we want the solution to satisfy some graph theoretic constraint also. For example, suppose a company wants to lease a set of mobile towers, each of which comes at a specific cost with a specific number of customers. The company has a budget and wants to serve the maximum number of customers. However, every mobile tower is in the range of some but not all the other towers. The company wants to provide fast connection between every pair of its customers, which is possible only if they are connected via other intermediary towers of the company. Here we have a natural graph structure — the vertices are the towers, and we have an edge between two vertices if they are in the range of one another. In this situation, the company wants to lease a connected subset of towers which has the maximum total number of customers (to maximize its earning) subject to its budget constraint. We call this problem Connected Knapsack.
Now, let us consider another application scenario where there is a railway network and a company wants to lease a railway path between station A and station B, maybe via other stations, to operate train services between them. Suppose the cost model puts a price tag on every station, and the links between two stations are complementary if the company owns both the stations. Each station also allows the owner to earn revenues from advertisement, etc. In this scenario, the company would like to lease from the entire railway network a path between stations A and B whose total revenues are as high as possible, subject to a budget constraint. We call this problem Path Knapsack. If the company’s primary goal is to provide the fastest connectivity between A and B, then the company wants to lease a shortest path between A and B, whose total revenues are as high as possible, subject to a budget constraint. We study this problem also under the name Shortest Path Knapsack. The formal definitions of the above problems are in Section 2.
1.1 Contribution
We study the computational complexity of Connected Knapsack, Path Knapsack, and Shortest Path Knapsack under the lens of parameterized complexity as well as approximation algorithms. We consider the treewidth of the underlying graph and the vertex cover size of the solution as our parameters. We note that both the parameters are well known to be small for many graphs appearing in various real-world applications. We summarize our results in Table 1.
Connected Knapsack Strongly $\mathsf{NP}$-complete even if max degree is $4$ [Theorem 3.1] $\mathsf{NP}$-complete even for stars [Theorem 3.2] $2^{\mathcal{O}(\text{tw}\log\text{tw})}\cdot\text{poly}(n)\min\{s^{2},d^{2}\}^{\dagger^{\star}}$ [Theorem 3.3] $2^{\mathcal{O}(\text{tw}\log\text{tw})}\cdot\text{poly}(n,1/\varepsilon)^{\dagger}$ time, $(1-\varepsilon)$ approximation [Theorem 3.4] No $\mathcal{O}(f(vcs).\text{poly}(n,s,d)$ algorithm unless $\mathsf{ETH}$ ${}^{\ddagger}$ [Theorem 3.5].
Path Knapsack Strongly $\mathsf{NP}$-complete even if max degree is $3$ [Theorem 4.1] $\mathsf{NP}$-complete for graphs with pathwidth 2 [Theorem 4.2] Polynomial-time algorithm for trees [Observation 1] $2^{\mathcal{O}(\text{tw}\log\text{tw})}\cdot\text{poly}(\min\{s^{2},d^{2}\})^{\dagger^{\star}}$ [Theorem 4.3] $2^{\mathcal{O}(\text{tw}\log\text{tw})}\cdot\text{poly}(n,1/\varepsilon)^{\dagger}$ time, $(1-\varepsilon)$ approximation [Theorem 4.4] $\mathcal{O}\left((2e)^{2vcs}vcs^{\mathcal{O}(\log vcs)}n^{\mathcal{O}(1)}\right)^{\ddagger}$ time algorithm [Corollary 2].
Shortest Path Knapsack $\mathsf{NP}$-complete for graphs with pathwidth 2 [Corollary 3] $\mathcal{O}(n\log n\cdot\min\{s^{2},(\alpha(\mathcal{V}))^{2}\})$ time algorithm [Theorem 5.1] Polynomial-time algorithm for trees [Observation 2] $(1-\varepsilon)$ approximation in $\text{poly}(n,1/\varepsilon)$ time [Theorem 5.2].
[…]
We observe that all our problems admit fixed-parameter pseudo-polynomial time algorithms with respect to the treewidth of the graph. Further, all the problems admit an FPTAS for graphs with small (at most $\mathcal{O}(\log n)$) treewidth. Our results seem to indicate that Shortest Path Knapsack is computationally the easiest, followed by Path Knapsack and Connected Knapsack.
1.2 Related Work
To the best of our knowledge, Yamada et al. [24] initiated the study of the Knapsack problem with a graph structure on the items (which they called conflict graph), where the goal is to compute an independent set which has the maximum total value and satisfies the budget constraint. The paper proposes a set of heuristics and upper bounds based on Lagrangean relaxation. Later, Hifi and Michrafy [15, 16] designed a meta-heuristic approach using reactive local search techniques for the problem. Pferschy and Schauer [22] showed that this problem is strongly $\mathsf{NP}$-hard and presented FPTAS for some special graph classes. Bettinelli et al. [2] proposed a dynamic programming based pruning in a branching based algorithm using the upper bounds proposed by Held et al. [14]. Coniglio et al. [7] presented a branch-and-bound type algorithm for the problem. Finally, Luiz et al. [19] proposed a cutting plane based algorithm. Pferschy and Schauer [23], Gurski and Rehs [13], and Goebbels et al. [12] showed approximation algorithms for the Knapsack problem with more sophisticated neighborhood constraints. This problem (and also the variant where the solution should be a clique instead of an independent set) also admits an algorithm running in time $n^{\mathcal{O}(k)}$ where $k$ is the thinness of the input graph [3, 20].
Ito et al. [17] studied FPTASes for a generalized version of our problems, called the Maximum partition problem but on a specialized graph class, namely the class of series-parallel graphs. Although their work mentions that their results can be extended to bounded treewidth graphs, the authors state that the algorithmic techniques for that is more complex but do not explain the techniques. To the best of our knowledge, there is no follow-up work where the techniques for bounded treewidth graphs are explained.
Bonomo-Braberman and Gonzalez [4] studied a general framework for, which they called "locally checkable problems", which can be used to obtain FPT algorithms for all our problems parameterized by treewidth. However, the running time of these algorithms obtained by their framework is worse than our algorithms parameterized by treewidth.
There are many other generalizations of the basic Knapsack problem which have been studied extensively in the literature. We refer to [5, 6, 18, 21] for an overview.
2 Preliminaries
We denote the set $\{0,1,2,\ldots\}$ of natural numbers with $\mathbb{N}$. For any integer $\ell$, we denote the sets $\{1,\ldots,\ell\}$ and $\{0,1,\ldots,\ell\}$ by $[\ell]$ and $[\ell]_{0}$ respectively. Given a graph $G=(V,E)$ the distance between two vertices $u,v\in V$ is denoted by ${\sf dist}_{G}(u,v)$. We now formally define our problems. We frame all our problems as decision problems so that we can directly use the framework of $\mathsf{NP}$-completeness.
Definition 1 (Connected Knapsack)
Given an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, non-negative weights of vertices $(w(u))_{u\in\mathcal{V}}$, non-negative values of vertices $(\alpha(u))_{u\in\mathcal{V}}$, size $s$ of the knapsack, and target value $d$, compute if there exists a subset $\mathcal{U}\subseteq\mathcal{V}$ of vertices such that: (i) $\mathcal{U}$ is connected, (ii) $w(\mathcal{U})=\sum_{u\in\mathcal{U}}w(u)\leqslant s$, and (iii) $\alpha(\mathcal{U})=\sum_{u\in\mathcal{U}}\alpha(u)\geqslant d$ We denote an arbitrary instance of Connected Knapsack by $(\mathcal{G},(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d)$.
If not mentioned otherwise, for any subset $\mathcal{V}^{\prime}\subseteq\mathcal{V}$, $\alpha(\mathcal{V}^{\prime})$ and $w(\mathcal{V}^{\prime})$, we denote $\sum_{u\in\mathcal{V}^{\prime}}\alpha(u)$ and $\sum_{u\in\mathcal{V}^{\prime}}w(u)$, respectively.
In our next problem, the knapsack subset of items must be a path between two given vertices.
Definition 2 (Path Knapsack)
Given an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, non-negative weights of vertices $(w(u))_{u\in\mathcal{V}}$, non-negative values of vertices $(\alpha(u))_{u\in\mathcal{V}}$, size $s$ of the knapsack, target value $d$, two vertices $x$ and $y$, compute if there exists a subset $\mathcal{U}\subseteq\mathcal{V}$ of vertices such that: (i) $\mathcal{U}$ is a path between $x$ and $y$ in $\mathcal{G}$, (ii) $w(\mathcal{U})=\sum_{u\in\mathcal{U}}w(u)\leqslant s$, and $\alpha(\mathcal{U})=\sum_{u\in\mathcal{U}}\alpha(u)\geqslant d$. We denote an arbitrary instance of Path Knapsack by $(\mathcal{G},(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d,x,y)$.
In Shortest Path Knapsack, the knapsack subset of items must be a shortest path between two given vertices.
Definition 3 (Shortest Path Knapsack)
Given an undirected edge-weighted graph $\mathcal{G}=(\mathcal{V},\mathcal{E},c:\mathcal{E}\longrightarrow\mathbb{R}^{+}_{\geqslant 0})$, non-negative weights of vertices $(w(u))_{u\in\mathcal{V}}$, non-negative values of vertices $(\alpha(u))_{u\in\mathcal{V}}$, size $s$ of knapsack, target value $d$, two vertices $x$ and $y$, compute if there exists a subset $\mathcal{U}\subseteq\mathcal{V}$ of vertices such that:
  1. 1.
    $\mathcal{U}$ is a shortest path between $x$ and $y$ in $\mathcal{G}$,
  2. 2.
    $w(\mathcal{U})=\sum_{u\in\mathcal{U}}w(u)\leqslant s$,
  3. 3.
    $\alpha(\mathcal{U})=\sum_{u\in\mathcal{U}}\alpha(u)\geqslant d$
We denote an arbitrary instance of Shortest Path Knapsack by $(\mathcal{G},(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d,x,y)$.
If not mentioned otherwise, we use $n,s,d,\mathcal{G}$ and $\mathcal{V}$ to denote respectively the number of vertices, the size of the knapsack, the target value, the input graph, and the set of vertices in the input graph.
Definition 4 (Treewidth)
Let $G=(V_{G},E_{G})$ be a graph. A tree-decomposition of a graph $G$ is a pair $(\mathbb{T}=(V_{\mathbb{T}},E_{\mathbb{T}}),\mathcal{X}=\{X_{t}\}_{t\in V_{\mathbb{T}}})$, where ${\mathbb{T}}$ is a tree where every node $t\in V_{\mathbb{T}}$ is assigned a subset $X_{t}\subseteq V_{G}$, called a bag, such that the following conditions hold.
[…]
    $\bigcup_{t\in V_{\mathbb{T}}}{X_{t}}=V_{G}$,
[…]
    for every edge $\{x,y\}\in E_{G}$ there is a $t\in V_{\mathbb{T}}$ such that $x,y\in X_{t}$, and
[…]
    for any $v\in V_{G}$ the subgraph of $\mathbb{T}$ induced by the set $\{t\mid v\in X_{t}\}$ is connected.
The width of a tree decomposition is $\max_{t\in V_{\mathbb{T}}}|X_{t}|-1$. The treewidth of $G$ is the minimum width over all tree decompositions of $G$ and is denoted by $\text{tw}(G)$.
A tree decomposition $(\mathbb{T},\mathcal{X})$ is called a nice edge tree decomposition if $\mathbb{T}$ is a tree rooted at some node $r$ where $X_{r}=\emptyset$, each node of $\mathbb{T}$ has at most two children, and each node is of one of the following kinds:
[…]
    Introduce node: a node $t$ that has only one child $t^{\prime}$ where $X_{t}\supset X_{t^{\prime}}$ and $|X_{t}|=|X_{t^{\prime}}|+1$.
[…]
    Introduce edge node a node $t$ labeled with an edge between $u$ and $v$, with only one child $t^{\prime}$ such that $\{u,v\}\subseteq X_{t^{\prime}}=X_{t}$. This bag is said to introduce $uv$.
[…]
    Forget vertex node: a node $t$ that has only one child $t^{\prime}$ where $X_{t}\subset X_{t^{\prime}}$ and $|X_{t}|=|X_{t^{\prime}}|-1$.
[…]
    Join node: a node $t$ with two children $t_{1}$ and $t_{2}$ such that $X_{t}=X_{t_{1}}=X_{t_{2}}$.
[…]
    Leaf node: a node $t$ that is a leaf of $\mathbb{T}$, and $X_{t}=\emptyset$.
We additionally require that every edge is introduced exactly once. One can show that a tree decomposition of width $t$ can be transformed into a nice tree decomposition of the same width $t$ and with $\mathcal{O}(t|V_{G}|)$ nodes, see e.g. [1, 9]. For a node $t\in\mathbb{T}$, let $\mathbb{T}_{t}$ be the subtree of $\mathbb{T}$ rooted at $t$, and $V(\mathbb{T}_{t})$ denote the vertex set in that subtree. Then $\beta(t)$ is the subgraph of $G$ where the vertex set is $\bigcup_{t^{\prime}\in V(\mathbb{T}_{t})}X_{t^{\prime}}$ and the edge set is the union of the set of edges introduced in each $t^{\prime},t^{\prime}\in V(\mathbb{T}_{t})$. We denote by $V(\beta(t))$ the set of vertices in that subgraph, and by $E(\beta(t))$ the set of edges of the subgraph.
In this paper, we sometimes fix a vertex $v\in V_{G}$ and include it in every bag of a nice edge tree decomposition $(\mathbb{T},\mathcal{X})$ of $G$, with the effect of the root bag and each leaf bag containing $v$. For the sake of brevity, we also call such a modified tree decomposition a nice tree decomposition. Given the tree $\mathbb{T}$ rooted at the node $r$, for any nodes $t_{1},t_{2}\in V_{\mathbb{T}}$, the distance between the two nodes in $\mathbb{T}$ is denoted by $\sf{dist}_{\mathbb{T}}(t_{1},t_{2})$.
3 Connected Knapsack
We present our results for Connected Knapsack in this section. First, we show that Connected Knapsack is strongly $\mathsf{NP}$-complete by reducing it from Vertex Cover, which is known to be $\mathsf{NP}$-complete even for $3$-regular graphs [10, folklore]. Hence, we do not expect a pseudo-polynomial time algorithm for Connected Knapsack, unlike Knapsack.
Definition 5 (Vertex Cover)
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a positive integer $k$, compute if there exists a subset $\mathcal{V}^{\prime}\subseteq\mathcal{V}$ such that at least one end-point of every edge belongs to $\mathcal{V}^{\prime}$ and $|\mathcal{V}^{\prime}|\leqslant k$. We denote an arbitrary instance of Vertex Cover by $(\mathcal{G},k)$.
Theorem 3.1
Connected Knapsack is strongly $\mathsf{NP}$-complete even when the maximum degree of the input graph is four.
Proof
Clearly, Connected Knapsack $\in$ $\mathsf{NP}$. We reduce Vertex Cover to Connected Knapsack to prove NP-hardness. Let $(\mathcal{G}=(\mathcal{V}=\{v_{i}:i\in[n]\},\mathcal{E}),$ $k)$ be an arbitrary instance of Vertex Cover where $\mathcal{G}$ is $3$-regular. We construct the following instance $(\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime}),(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d)$ of Connected Knapsack.
$\mathcal{V}^{\prime}=\{u_{i},g_{i}:i\in[n]\}\cup\{h_{e}:e\in\mathcal{E}\}$
=$\mathcal{E}^{\prime}=\{\{u_{i},h_{e}\}:i\in[n],e\in\mathcal{E},e\text{ is incident on }v_{i}\text{ in }\mathcal{G}\}$
$\cup\{\{u_{i},g_{i}\}:i\in[n]\}\cup\{\{g_{i},g_{i+1}\}:i\in[n-1]\}$
$w(u_{i})=1,\alpha(u_{i})=0,w(g_{i})=0,\alpha(g_{i})=0,\forall i\in[n]$
$w(h_{e})=0,\alpha(h_{e})=1,\forall e\in\mathcal{E},s=k,d=|\mathcal{E}|$
We observe that the maximum degree of $\mathcal{G}^{\prime}$ is at most four — (i) the degree of $u_{i}$ is four for every $i\in[n]$, since $\mathcal{G}$ is $3$-regular and $u_{i}$ has an edge to $g_{i}$, (ii) the degree of $h_{e}$ is two for every $e\in\mathcal{E}$, and (iii) the degree of $g_{i}$ is at most three for every $i\in[n]$ since the set $\{g_{i}:i\in[n]\}$ induces a path. We claim that the two instances are equivalent.
In one direction, let us suppose that the Vertex Cover instance is a yes instance. Let $\mathcal{W}\subseteq\mathcal{V}$ be a vertex cover of $\mathcal{G}$ with $|\mathcal{W}|\leqslant k$. We consider $\mathcal{U}=\{u_{i}:i\in[n],v_{i}\in\mathcal{W}\}\cup\{g_{i}:i\in[n]\}\cup\{h_{e}:e\in\mathcal{E}\}\subseteq\mathcal{V}^{\prime}$. We claim that $\mathcal{G}^{\prime}[\mathcal{U}]$ is connected. Since $\{g_{i}:i\in[n]\}$ induces a path and there is an edge between $u_{i}$ and $g_{i}$ for every $i\in[n]$, the induced subgraph $\mathcal{G}^{\prime}[\{u_{i}:i\in[n],v_{i}\in\mathcal{W}\}\cup\{g_{i}:i\in[n]\}]$ is connected. Since $\mathcal{W}$ is a vertex cover of $\mathcal{G}$, every edge $e\in\mathcal{E}$ is incident on at least one vertex in $\mathcal{W}$. Hence, every vertex $h_{e},e\in\mathcal{E},$ has an edge with at least one vertex in $\{u_{i}:i\in[n],v_{i}\in\mathcal{W}\}$ in the graph $\mathcal{G}^{\prime}$. Hence, the induced subgraph $\mathcal{G}^{\prime}[\mathcal{U}]$ is connected. Now we have $w(\mathcal{U})=\sum_{i=1}^{n}w(u_{i})\mathbbm{1}(u_{i}\in\mathcal{U})+\sum_{i=1}^{n}w(g_{i})+\sum_{e\in\mathcal{E}}w(h_{e})=|\mathcal{U}|\leqslant k$. We also have $\alpha(\mathcal{U})=\sum_{i=1}^{n}\alpha(u_{i})\mathbbm{1}(u_{i}\in\mathcal{U})+\sum_{i=1}^{n}\alpha(g_{i})+\sum_{e\in\mathcal{E}}\alpha(h_{e})=|\mathcal{E}|$. Hence, the Connected Knapsack instance is also a yes instance.
In the other direction, let us assume that the Connected Knapsack instance is a yes instance. Let $\mathcal{U}\subseteq\mathcal{V}^{\prime}$ be a solution of the Connected Knapsack instance. We consider $\mathcal{W}=\{v_{i}:i\in[n]:u_{i}\in\mathcal{U}\}$. Since $s=k$, we have $|\mathcal{W}|\leqslant k$. Also, since $d=|\mathcal{E}|$, we have $\{h_{e}:e\in\mathcal{E}\}\subseteq\mathcal{U}$. We claim that $\mathcal{W}$ is a vertex cover of $\mathcal{G}$. Suppose not, then there exists an edge $e\in\mathcal{E}$ which is not covered by $\mathcal{W}$. Then none of the neighbors of $h_{e}$ belongs to $\mathcal{U}$ contradicting our assumption that $\mathcal{U}$ is a solution and thus $\mathcal{G}^{\prime}[\mathcal{U}]$ should be connected. Hence, $\mathcal{W}$ is a vertex cover of $\mathcal{G}$ and thus the Vertex Cover instance is a yes instance.
We observe that all the numbers in our reduced Connected Knapsack instance are at most the number of edges of the graph. Hence, our reduction shows that Connected Knapsack is strongly $\mathsf{NP}$-complete.
Theorem 3.1 also implies the following corollary in the framework of parameterized complexity.
Corollary 1
[…]
We next show that Connected Knapsack is $\mathsf{NP}$-complete even for trees. For that, we reduce from the $\mathsf{NP}$-complete problem Knapsack.
Definition 6 (Knapsack)
Given a set $\mathcal{X}=[n]$ of $n$ items with sizes $\theta_{1},\ldots,\theta_{n},$ values $p_{1},\ldots,p_{n}$, capacity $b$ and target value $q$, compute if there exists a subset $\mathcal{I}\subseteq[n]$ such that $\sum_{i\in\mathcal{I}}\theta_{i}\leqslant b$ and $\sum_{i\in\mathcal{I}}p_{i}\geqslant q$. We denote an arbitrary instance of Knapsack by $(\mathcal{X},(\theta_{i})_{i\in\mathcal{X}},(p_{i})_{i\in\mathcal{X}},b,q)$.
Theorem 3.2
Connected Knapsack is $\mathsf{NP}$-complete even for star graphs.
Proof
Connected Knapsack clearly belongs to $\mathsf{NP}$. To show $\mathsf{NP}$-hardness, we reduce from Knapsack. Let $(\mathcal{X}=[n],(\theta_{i})_{i\in\mathcal{X}},(p_{i})_{i\in\mathcal{X}},b,q)$ be an arbitrary instance of Knapsack. We consider the following instance $(\mathcal{G}(\mathcal{V},\mathcal{E}),(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d)$ of Connected Knapsack.
$\mathcal{V}=\{v_{0},v_{1},\ldots,v_{n}\}$
=$\mathcal{E}=\{\{v_{0},v_{i}\}:1\leqslant i\leqslant n\}$
$w(v_{i})=\theta_{i}\,\alpha(v_{i})=p_{i};\forall i\in[n],w(v_{0})=\alpha(v_{0})=0;$
$s=b,d=q$
We now claim that the two instances are equivalent.
In one direction, let us suppose that the Knapsack instance is a yes instance. Let $\mathcal{W}\subseteq\mathcal{X}$ be a solution of Knapsack. Let us consider $\mathcal{U}=\{v_{i}:i\in\mathcal{W}\}\cup\{v_{0}\}\subseteq\mathcal{V}$. We observe that $\mathcal{G}[\mathcal{U}]$ is connected since $v_{0}\in\mathcal{U}$. We also have
$w(\mathcal{U})=\sum_{i\in\mathcal{W}}w(v_{i})=\sum_{i\in\mathcal{W}}\theta_{i}\leqslant b=s,$
and
$\alpha(\mathcal{U})=\sum_{i\in\mathcal{W}}\alpha(v_{i})=\sum_{i\in\mathcal{W}}p_{i}\geqslant q=d.$
Hence, the Connected Knapsack instance is a yes instance.
In the other direction, let us assume that the Connected Knapsack instance is a yes instance with $\mathcal{U}\subseteq\mathcal{V}$ be one of its solution. Let us consider a set $\mathcal{W}=\{i:i\in[n],v_{i}\in\mathcal{U}\}$. We now have
$\sum_{i\in\mathcal{W}}\theta_{i}=\sum_{i\in\mathcal{W}}w(v_{i})=w(\mathcal{U})\leqslant s=b,$
and
$\sum_{i\in\mathcal{W}}p_{i}=\sum_{i\in\mathcal{W}}\alpha(v_{i})=\alpha(\mathcal{U})\geqslant d=q.$
Hence, the Knapsack instance is a yes instance.
We complement the hardness result in Theorem 3.2 by designing a pseudo-polynomial-time algorithm for Connected Knapsack for trees. In fact, we have designed an algorithm with running time $2^{\mathcal{O}(\text{tw}\log\text{tw})}\cdot n^{\mathcal{O}(1)}\cdot{\sf min}\{s^{2},d^{2}\}$ where the treewidth of the input graph is tw. We present this algorithm next.
3.1 Treewidth as a Parameter
Theorem 3.3
[…]
Proof
Let $(G=(V_{G},E_{G}),{(w(u))_{u\in V_{G}},(\alpha(u))_{u\in V_{G}}},s,d)$ be an input instance of Connected Knapsack such that $\text{tw}=tw(G)$. Let $\mathcal{U}\subseteq V_{G}$ be a solution subset for the input instance. For technical purposes, we guess a vertex $v\in\mathcal{U}$ — once the guess is fixed, we are only interested in finding solution subsets $\mathcal{U}^{\prime}$ that contain $v$ and $\mathcal{U}$ is one such candidate. We also consider a nice edge tree decomposition $(\mathbb{T}=(V_{\mathbb{T}},E_{\mathbb{T}}),\mathcal{X})$ of $G$ that is rooted at a node $r$, and where $v$ has been added to all bags of the decomposition. Therefore, $X_{r}=\{v\}$ and each leaf bag is the singleton set $\{v\}$.
We define a function $\ell:V_{\mathbb{T}}\rightarrow\mathbb{N}$. For a vertex $t\in V_{\mathbb{T}}$, $\ell(t)=\sf{dist}_{\mathbb{T}}(t,r)$, where $r$ is the root. Note that this implies that $\ell(r)=0$. Let us assume that the values that $\ell$ takes over the nodes of $\mathbb{T}$ is between 00 and $L$. Now, we describe a dynamic programming algorithm over $(\mathbb{T},\mathcal{X})$ for Connected Knapsack.
States.
We maintain a DP table $D$ where a state has the following components:
  1. 1.
    $t$ represents a node in $V_{\mathbb{T}}$.
  2. 2.
    $\mathbb{P}=(P_{0},\ldots,P_{m}),m\leqslant\text{tw}+1$ represents a partition of the vertex subset $X_{t}$.
Interpretation of States.
For a state $[t,\mathbb{P}]$, if there is a solution subset $\mathcal{U}$ let $\mathcal{U^{\prime}}=\mathcal{U}\cap V(\beta(t))$. Let $\beta_{\mathcal{U^{\prime}}}$ be the graph induced on $\mathcal{U^{\prime}}$ in $\beta(t)$. Let $C_{1},C_{2},\ldots,C_{m}$ be the connected components of $\beta_{\mathcal{U^{\prime}}}$. Note that $m\leqslant\text{tw}+1$. Then in the partition $\mathbb{P}=(P_{0},P_{1},\ldots,P_{m})$, $P_{i}=C_{i}\cap X_{t},1\leqslant i\leqslant m$. Also, $P_{0}=X_{t}\setminus\mathcal{U^{\prime}}$.
Given a node $t\in V_{\mathbb{T}}$, a subgraph $H$ of $\beta(t)$ is said to be a $\mathbb{P}$-subgraph if (i) the connected components $C_{1},C_{2},\ldots,C_{m}$ of $H$, $m\leqslant\text{tw}$ are such that $P_{i}=C_{i}\cap X_{t},1\leqslant i\leqslant m$, (ii) $P_{0}=X_{t}\setminus H$. For each state $[t,\mathbb{P}]$, a pair $(w,\alpha)$ with $w\leqslant s$ is said to be feasible if there is a $\mathbb{P}$-subgraph of $\beta(t)$ whose total weight is $w$ and total value is $\alpha$. Moreover, a feasible pair $(w,\alpha)$ is said to be undominated if there is no other $\mathbb{P}$-subgraph with weight $w^{\prime}$ and value $\alpha^{\prime}$ such that $w^{\prime}\leqslant w$ and $\alpha^{\prime}\geqslant\alpha$. Please note that by default, an empty $\mathcal{P}$-graph has total weight 00 and total value 00.
For each state $[t,\mathbb{P}]$, we initialize $D[t,\mathbb{P}]$ to the list $\{(0,0)\}$. Our computation shall be such that in the end each $D[t,\mathbb{P}]$ stores the set of all undominated feasible pairs $(w,\alpha)$ for the state $[t,\mathbb{P}]$.
Dynamic Programming on $D$.
We describe the following procedure to update the table $D$. We start updating the table for states with nodes $t\in V_{\mathbb{T}}$ such that $\ell(t)=L$. When all such states are updated, then we move to update states where the node $t$ has $\ell(t)=L-1$, and so on till we finally update states with $r$ as the node — note that $\ell(r)=0$. For a particular $i$, $0\leqslant i<L$ and a state $[t,\mathbb{P}]$ such that $\ell(t)=i$, we can assume that $D[t^{\prime},\mathbb{P}^{\prime}]$ have been evaluated for all $t^{\prime}$ such that $\ell(t^{\prime})>i$ and all partitions $\mathbb{P}^{\prime}$ of $X_{t^{\prime}}$. Now we consider several cases by which $D[t,\mathbb{P}]$ is updated based on the nature of $t$ in $\mathbb{T}$:
  1. 1.
    Suppose $t$ is a leaf node. Note that by our modification, $X_{t}=\{v\}$. There can be only 2 partitions for this singleton set — $\mathbb{P}_{t}^{1}=(P_{0}=\emptyset,P_{1}=\{v\})$ and $\mathbb{P}_{t}^{2}=(P_{0}=\{v\},P_{1}=\emptyset)$. If $\mathbb{P}=\mathbb{P}_{t}^{1}$ then $D[t,\mathbb{P}]$ stores the pair $(w(v),\alpha(v))$ if $w(v)\leqslant s$ and otherwise no modification is made. If $\mathbb{P}=\mathbb{P}_{t}^{2}$ then $D[t,\mathbb{P}]$ is not modified.
  2. 2.
    Suppose $t$ is a forget vertex node. Then it has an only child $t^{\prime}$ where $X_{t}\subset X_{t^{\prime}}$ and there is exactly one vertex $u\neq v$ that belongs to $X_{t^{\prime}}$ but not to $X_{t}$. Let $\mathbb{P}^{\prime}=(P^{\prime}_{0},P^{\prime}_{1},\ldots,P^{\prime}_{m^{\prime}})$ be a partition of $X_{t^{\prime}}$ such that when restricted to $X_{t}$ we obtain the partition $\mathbb{P}=(P_{0},P_{1},\ldots,P_{m})$. For each such partition, we shall do the following.
    Suppose $\mathbb{P}^{\prime}$ has $u\in P^{\prime}_{0}$, then each feasible undominated pair stored in $D[t^{\prime},\mathbb{P}^{\prime}]$ is copied to $D[t,\mathbb{P}]$.
    Alternatively, suppose $\mathbb{P}^{\prime}$ has $u\in P^{\prime}_{i},i>0$ and $|P^{\prime}_{i}|>1$. Then, each feasible undominated pair stored in $D[t^{\prime},\mathbb{P}^{\prime}]$ is copied to $D[t,\mathbb{P}]$.
    Finally, suppose $\mathbb{P}^{\prime}$ has $u\in P^{\prime}_{i},i>0$ and $P^{\prime}_{i}=\{u\}$. Then we do not make any changes to $D[t,\mathbb{P}]$.
  3. 3.
    Suppose $t$ is an introduce node. Then it has an only child $t^{\prime}$ where $X_{t^{\prime}}\subset X_{t}$ and there is exactly one vertex $u\neq v$ that belongs to $X_{t}$ but not $X_{t^{\prime}}$. Note that no edges incident to $u$ have been introduced yet, and so in $\beta(t)$ $u$ is not yet connected to any other vertex. Let $\mathbb{P}^{\prime}=(P^{\prime}_{0},P^{\prime}_{1},\ldots,P^{\prime}_{m^{\prime}})$ be the partition of $X_{t^{\prime}}$ obtained from restricting the partition $\mathbb{P}=(P_{0},P_{1},\ldots,P_{m})$ to $X_{t^{\prime}}$. First, suppose $u\in P_{0}$. Then we copy all pairs of $D[t^{\prime},\mathbb{P}^{\prime}]$ to $D[t,\mathbb{P}]$.
    Next, suppose $u\in P_{i},i>0$ and $P_{i}=\{u\}$. Then for each pair $(w,\alpha)$ in $D[t^{\prime},\mathbb{P}^{\prime}]$, if $w+w(u)\leqslant s$ we add $(w+w(u),\alpha+\alpha(u))$ to the set in $D[t,\mathbb{P}]$.
    Finally, let $u\in P_{i},i>0$ and $|P_{i}|>1$. Then we make no changes to $D[t,\mathbb{P}]$.
  4. 4.
    Suppose $t$ is an introduce edge node. Then it has an only child $t^{\prime}$ where $X_{t^{\prime}}=X_{t}$, and additionally for two vertices $u,w\in X_{t}=X_{t^{\prime}}$, the edge $\{u,w\}$ is introduced into $\beta(t)$. First, suppose $\mathbb{P}=(P_{0},P_{1},\ldots,P_{m})$ is such that one of $u,w$ is in $P_{0}$. Then all pairs of $D[t^{\prime},\mathbb{P}]$ are copied to $D[t,\mathbb{P}]$.
    Next, let $u\in P_{i}$, $w\in P_{j}$, $i\neq j\neq 0$. Then no updates are made to $D[t,\mathbb{P}]$.
    Finally, suppose $u,w\in P_{i},i>0$. Copy to $D[t,\mathbb{P}]$ all pairs from $D[t^{\prime},\mathbb{P}]$. Consider a partition $\mathbb{P}^{\prime}$ where the part $P^{\prime}$ contains $u$ and $P^{\prime\prime}$ contains $w$. $\mathbb{P}^{\prime}$ is such that $P_{i}=P^{\prime}\cup P^{\prime\prime}$ and any other $P_{j}$ with $j\neq i$ is a part in $\mathbb{P}^{\prime}$. Any pair of $D[t^{\prime},\mathbb{P}^{\prime}]$ is copied to $D[t,\mathbb{P}]$.
  5. 5.
    Suppose $t$ is a join node. Then it has two children $t_{1},t_{2}$ such that $X_{t}=X_{t_{1}}=X_{t_{2}}$. Consider $\mathbb{P}=(P_{0},P_{1},\ldots,P_{m})$. Let $(w_{\mathcal{P}},\alpha_{\mathcal{P}})$ be the total weight and value of the vertices in $\cup_{1\leqslant i\leqslant m}P_{i}$. Consider a pair $(w_{1},\alpha_{1})$ in $D[t_{1},\mathbb{P}]$ and a pair $(w_{2},\alpha_{2})$ in $D[t_{2},\mathbb{P}]$. Suppose $w_{1}+w_{2}-w_{\mathbb{P}}\leqslant s$. Then we add $(w_{1}+w_{2}-w_{\mathbb{P}},\alpha_{1}+\alpha_{2}-\alpha_{\mathbb{P}})$ to $D[t,\mathbb{P}]$.
Finally, in the last step of updating $D[t,\mathbb{P}]$, we go through the list saved in $D[t,\mathbb{P}]$ and only keep undominated pairs.
The output of the algorithm is a pair $(w,\alpha)$ stored in $D[r,\mathbb{P}=(P_{0}=\emptyset,P_{1}=\{v\})]$ such that $w\leqslant s$ and $\alpha$ is the maximum value over all pairs in $D[r,\mathbb{P}]$.
Correctness of the Algorithm.
Recall that we are looking for a solution $\mathcal{U}$ that contains the fixed vertex $v$ that belongs to all bags of the tree decomposition. First, we show that a pair $(w,\alpha)$ belonging to $D[t,\mathbb{P}]$ for a node $t\in V_{\mathbb{T}}$ and a partition $\mathbb{P}$ of $X_{t}$ corresponds to a $\mathbb{P}$-subgraph $H$ in $\beta(t)$. Recall that $X_{r}=\{v\}$. Thus, this implies that a pair $(w,\alpha)$ belonging to $D[r,\mathbb{P}=(P_{0}=\emptyset,P_{1}=\{v\})]$ corresponds to a connected subgraph of $G$. Moreover, the output is a pair that is feasible and with the highest value.
In order to show that a pair $(w,\alpha)$ belonging to $D[t,\mathbb{P}]$ for a node $t\in V_{\mathbb{T}}$ and a partition $\mathbb{P}$ of $X_{t}$ corresponds to a $\mathbb{P}$-subgraph $H$ in $\beta(t)$, we need to consider the cases of what $t$ can be:
  1. 1.
    When $t$ is a leaf node with $\ell(t)=i$, $X_{t}$ only contains $v$ and the update to $D$ is done such that $v$ is the corresponding subgraph to a stored pair. This is true in particular when $i=L$, the base case. From now we can assume that for a node $t$ with $\ell(t)=i<L$ all $D[t^{\prime},\mathbb{P}^{\prime}]$ entries are correct and correspond to $\mathbb{P}^{\prime}$-subgraphs in $\beta(t^{\prime})$ when $\ell(t^{\prime})>i$.
  2. 2.
    When $t$ is a forget vertex node, let $t^{\prime}$ be the child node and $u\neq v$ be the vertex that is being forgotten. We copy pairs from $D[t^{\prime},\mathbb{P}^{\prime}]$ depending on the structure of $\mathbb{P}^{\prime}$. Since $\ell(t^{\prime})>\ell(t)$, by induction hypothesis all entries in $D[t^{\prime},\mathbb{P}^{\prime}]$ for any partition $\mathbb{P}^{\prime}$ of $X_{t^{\prime}}$ are feasible. From the cases considered, we copy a pair to $D[t,\mathbb{P}]$ from a $D[t^{\prime},\mathbb{P}^{\prime}]$ only when $u$ is not part of the $\mathbb{P}^{\prime}$-subgraph or is in a component of $\mathbb{P}^{\prime}$ that has vertices in $X_{t}$. Thus, the same subgraph is a $\mathbb{P}$-subgraph in $\beta(t)$.
  3. 3.
    When $t$ is an introduce node, there is a child $t^{\prime}$ we are introducing a vertex $u\neq v$ that has no adjacent edges added in $\beta(t)$. Since $\ell(t^{\prime})>\ell(t)$, by induction hypothesis all entries in $D[t^{\prime},\mathbb{P}^{\prime}]$ for any partition $\mathbb{P}^{\prime}$ of $X_{t^{\prime}}$ are feasible. We update pairs in $D[t,\mathbb{P}]$ from $D[t^{\prime},\mathbb{P}^{\prime}]$ such that either $u$ is not considered as part of a $\mathbb{P}$-subgraph and the pair is certified by the $\mathbb{P}^{\prime}$-subgraph, or $u$ is added to a $\mathbb{P}^{\prime}$-subgraph in order to obtain a new $\mathbb{P}$-subgraph.
  4. 4.
    When $t$ is an introduce edge node, there is a child $t^{\prime}$ such that $X_{t}=X_{t^{\prime}}$ and the only difference is that two vertices $u,w$ in the bags $X_{t}=X_{t^{\prime}}$ now have an edge in $\beta(t)$. Since $\ell(t^{\prime})>\ell(t)$, by induction hypothesis all entries in $D[t^{\prime},\mathbb{P}^{\prime}]$ for any partition $\mathbb{P}^{\prime}$ of $X_{t^{\prime}}$ are feasible. The updates are made in the cases when one of $u$ or $w$ is not in the intended $\mathbb{P}$-subgraph and the included pair is certified by a $\mathbb{P^{\prime}}$-subgraph, or when the $u$ and $w$ belong to different components of a $\mathbb{P}^{\prime}$-subgraph and the new $\mathbb{P}$-subgraph has these components merged as a single component.
  5. 5.
    When $t$ is a join node, there are two children $t_{1},t_{2}$ such that $X_{T}=X_{t_{1}}=X_{t_{2}}$. Note that this implies that $V(\beta(t_{1}))\cap V(\beta(t_{2}))=X_{t}$. Since $\ell(t_{1}),\ell(t_{2})>\ell(t)$, by induction hypothesis all entries in $D[t_{i},\mathbb{P}^{\prime}]$ for any partition $\mathbb{P}^{\prime}$ of $X_{t_{i}}$ are feasible for $i\in\{1,2\}$.We update pairs in $D[t,\mathbb{P}]$ when there is a $\mathbb{P}$-subgraph in $\beta(t_{1})$ and a $\mathbb{P}$-subgraph in $\beta(t_{2})$ and we take the union of these two subgraphs to obtain a $\mathbb{P}$-subgraph in $\beta(t)$.
Thus in all cases of $t$, a pair added to $D[t,\mathbb{P}]$ for some partition of $X_{t}$ is a feasible pair. Recall that as a last step of updating $D[t,\mathbb{P}]$, we go through the entire list and keep only undominated pairs in the list.
What remains to be shown is that an undominated feasible solution $\mathcal{U}$ of Connected Knapsack in $G$ is contained in $D[r,\mathbb{P}=(P_{0}=\emptyset,P_{1}=\{v\})]$. Let $w$ be the weight of $\mathcal{U}$ and $\alpha$ be the value. Recall that $v\in\mathcal{U}$. For each $t$, we consider the subgraph $\beta(t)\cap\mathcal{U}$. Let $C_{1},C_{2},\ldots,C_{m}$ be components of $\beta(t)\cap\mathcal{U}$ and let for each $1\leqslant i\leqslant m,P_{i}=X_{t}\cap C_{i}$. Also, let $P_{0}=X_{t}\setminus\mathcal{U}$. Consider $\mathbb{P}=(P_{0},P_{1},\ldots,P_{m})$. The algorithm updates in $D[t,\mathbb{P}]$ the pair $(w^{\prime},\alpha^{\prime})$ for the subsolution $\beta(t)\cap\mathcal{U}$. Therefore, $D[r,\mathbb{P}=(\emptyset,\{v\})]$ contains the pair $(w,\alpha)$. Thus, we are done.
Running time.
There are $n$ choices for the fixed vertex $v$. Upon fixing $v$ and adding it to each bag of $(\mathbb{T},\mathcal{X})$ we consider the total possible number of states. There are at most $\mathcal{O}(n)\cdot 2^{\text{tw}\log\text{tw}}$ states. For each state, since we are keeping only undominated pairs, for each $w$ there can be at most one pair with $w$ as the first coordinate; similarly, for each $\alpha$ there can be at most one pair with $\alpha$ as the second coordinate. Thus, the number of undominated pairs in each $D[t,\mathbb{P}]$ is at most ${\sf min}\{s,d\}$. By the description of the algorithm, the maximum length of the list stored at $D[t,\mathbb{P}]$ during updation, but before the check is made for only undominated pairs, is $2^{\text{tw}\log\text{tw}}\cdot{\sf min}\{s^{2},d^{2}\}$. Thus, updating the DP table at any vertex takes $2^{\text{tw}\log\text{tw}}\cdot{\sf min}\{s^{2},d^{2}\}$ time. Since there are $\mathcal{O}(n)$ vertices in $\mathbb{T}$, the total running time of the algorithm is $2^{\mathcal{O}(\text{tw}\log\text{tw})}\cdot n\cdot{\sf min}\{s^{2},d^{2}\}$.
3.2 A Fixed Parameter Fully Pseudo-polynomial Time Approximation Scheme
We now use the algorithm in Theorem 3.3 as a black-box to design an $(1-\varepsilon)$-factor approximation algorithm for optimizing the value of the solution and running in time $2^{\text{tw}\log\text{tw}}\cdot\text{poly}(n,1/\varepsilon)$.
Theorem 3.4
[…]
Proof
Let $\mathcal{I}=(\mathcal{G}=(\mathcal{V},\mathcal{E}),(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s)$ be an arbitrary instance of Connected Knapsack where the goal is to output a connected subgraph $\mathcal{U}$ of maximum $\alpha(\mathcal{U})$ subject to the constraint that $w(\mathcal{U})\leqslant s$. Without loss of generality, we can assume that $w(u)\leqslant s$ for every $u\in\mathcal{V}$. If not, then we can remove every $u\in\mathcal{V}$ whose $w(u)>s$; this does not affect any solution since any vertex deleted can never be part of any solution. Let $\alpha_{\text{max}}=\max\{\alpha(u):u\in\mathcal{V}\}$. We construct another instance $\mathcal{I}^{\prime}=\left(\mathcal{G}=(\mathcal{V},\mathcal{E}),(w(u))_{u\in\mathcal{V}},(\alpha^{\prime}(u)=\left\lfloor\frac{n\alpha(u)}{\varepsilon\alpha_{\text{max}}}\right\rfloor)_{u\in\mathcal{V}},s\right)$ of Connected Knapsack. We compute the optimal solution $\mathcal{W}^{\prime}\subseteq\mathcal{V}$ of $\mathcal{I}^{\prime}$ using the algorithm in Theorem 3.3 and output $\mathcal{W}^{\prime}$. Let $\mathcal{W}\subseteq\mathcal{V}$ be an optimal solution of $\mathcal{I}$. Clearly $\mathcal{W}^{\prime}$ is a valid (may not be optimal) solution of $\mathcal{I}$ also, since $w(\mathcal{W}^{\prime})\leqslant s$ by the correctness of the algorithm in Theorem 3.3. We now prove the approximation factor of our algorithm.
$\sum_{u\in\mathcal{W}^{\prime}}\alpha(u)$ $\geqslant\frac{\varepsilon\alpha_{\text{max}}}{n}\sum_{u\in\mathcal{W}^{\prime}}\left\lfloor\frac{n\alpha(u)}{\varepsilon\alpha_{\text{max}}}\right\rfloor$
$\geqslant\frac{\varepsilon\alpha_{\text{max}}}{n}\sum_{u\in\mathcal{W}}\left\lfloor\frac{n\alpha(u)}{\varepsilon\alpha_{\text{max}}}\right\rfloor$ [since $\mathcal{W}^{\prime}$ is an optimal solution of $\mathcal{I}^{\prime}$]
$\geqslant\frac{\varepsilon\alpha_{\text{max}}}{n}\sum_{u\in\mathcal{W}}\left(\frac{n\alpha(u)}{\varepsilon\alpha_{\text{max}}}-1\right)$
$\geqslant\left(\sum_{u\in\mathcal{W}}\alpha(u)\right)-\varepsilon\alpha_{\text{max}}$
$\geqslant\mathsf{OPT}(\mathcal{I})-\varepsilon\mathsf{OPT}(\mathcal{I})$ [$\alpha_{\text{max}}\leqslant\mathsf{OPT}(\mathcal{I})$]
$=(1-\varepsilon)\mathsf{OPT}(\mathcal{I})$
Hence, the approximation factor of our algorithm is $(1-\varepsilon)$. We now analyze the running time of our algorithm.
The value of any optimal solution of $\mathcal{I}^{\prime}$ is at most
$\sum_{u\in\mathcal{V}}\alpha^{\prime}(u)\leqslant\frac{n}{\varepsilon\alpha_{\text{max}}}\sum_{u\in\mathcal{V}}\alpha(u)\leqslant\frac{n}{\varepsilon\alpha_{\text{max}}}\sum_{u\in\mathcal{V}}\alpha_{\text{max}}=\frac{n^{2}}{\varepsilon}.$
Hence, the running time of our algorithm is $2^{\text{tw}\log\text{tw}}\cdot\text{poly}(n,1/\varepsilon)$.
3.2.1 Other Parameters
[…]
Definition 7 (Partial Vertex Cover)
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and two integers $k$ and $\ell$, compute if there exists a subset $\mathcal{V}^{\prime}\subseteq\mathcal{V}$ such that (i) $|\mathcal{V}^{\prime}|\leqslant k$ and (ii) there exist at least $\ell$ edges whose one or both end points belong to $\mathcal{V}^{\prime}$. We denote an arbitrary instance of Partial Vertex Cover by $(\mathcal{G},k,\ell)$.
Theorem 3.5
[…]
Proof
At a high-level, our reduction from Partial Vertex Cover is similar to the reduction in Theorem 3.1. The only difference is that we will now have only one “global vertex” instead of the “global path” in the proof of Theorem 3.1. Formally, our reduction is as follows.
Let $(\mathcal{G}=(\mathcal{V},\mathcal{E}),k,\ell)$ be an arbitrary instance of Partial Vertex Cover. We consider the following instance $(\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime}),(w(u))_{u\in\mathcal{V}^{\prime}},(\alpha(u))_{u\in\mathcal{V}^{\prime}},s,d)$ of Connected Knapsack.
$\mathcal{V}^{\prime}=\{u_{i}:i\in[n]\}\cup\{h_{e}:e\in\mathcal{E}\}\cup\{g\}$
=$\mathcal{E}^{\prime}=\{\{u_{i},h_{e}\}:i\in[n],e\in\mathcal{E},e\text{ is incident on }v_{i}\text{ in }\mathcal{G}\}\cup\{\{u_{i},g\}:i\in[n]\}$
$w(u_{i})=1,\alpha(u_{i})=0,\forall i\in[n],w(h_{e})=0,\alpha(h_{e})=1,\forall e\in\mathcal{E},$
$w(g)=0,\alpha(g)=0,s=k,d=\ell$
We claim that the two instances are equivalent.
In one direction, let us suppose that the Partial Vertex Cover instance is a yes instance. Let $\mathcal{W}\subseteq\mathcal{V}$ covers at least $\ell$ edges in $\mathcal{G}$ and $|\mathcal{W}|\leqslant k$. We consider $\mathcal{U}=\{u_{i}:i\in[n],v_{i}\in\mathcal{W}\}\cup\{h_{e}:e\in\mathcal{E},e\text{ is covered by }\mathcal{W}\}\cup\{g\}\subseteq\mathcal{V}^{\prime}$. We claim that $\mathcal{G}^{\prime}[\mathcal{U}]$ is connected. Since there is an edge between $u_{i}$ and $g$ for every $i\in[n]$, the induced subgraph $\mathcal{G}^{\prime}[\{u_{i}:i\in[n],v_{i}\in\mathcal{W}\}\cup\{g\}]$ is connected. Also, since $h_{e}$ belongs to $\mathcal{U}$ only if $\mathcal{W}$ covers $e$ in $\mathcal{G}$, the induced subgraph $\mathcal{G}^{\prime}[\mathcal{U}]$ is connected. Now we have $w(\mathcal{U})=\sum_{i=1}^{n}w(u_{i})\mathbbm{1}(u_{i}\in\mathcal{U})+w(g)+\sum_{e\in\mathcal{E},\mathcal{W}\text{ covers }e}w(h_{e})=|\mathcal{W}|\leqslant k$. We also have $\alpha(\mathcal{U})=\sum_{i=1}^{n}\alpha(u_{i})\mathbbm{1}(u_{i}\in\mathcal{U})+\alpha(g)+\sum_{e\in\mathcal{E},\mathcal{W}\text{ covers }e}\alpha(h_{e})\geqslant\ell$. Hence, the Connected Knapsack instance is also a yes instance.
In the other direction, let us assume that the Connected Knapsack instance is a yes instance. Let $\mathcal{U}\subseteq\mathcal{V}^{\prime}$ be a solution of the Connected Knapsack instance. We consider $\mathcal{W}=\{v_{i}:i\in[n]:u_{i}\in\mathcal{U}\}$. Since $s=k$, we have $|\mathcal{W}|\leqslant k$. Also, since $d=\ell$, we have |$|\{h_{e}:e\in\mathcal{E}\}\cap\mathcal{U}|\geqslant\ell$. We claim that $\mathcal{W}$ covers every edge in $\{e:e\in\mathcal{E},h_{e}\in\mathcal{U}\}$. Suppose not, then there exists an edge $e\in\{e:e\in\mathcal{E},h_{e}\in\mathcal{U}\}$ which is not covered by $\mathcal{W}$. Then none of the neighbors of $h_{e}$ belongs to $\mathcal{U}$ contradicting our assumption that $\mathcal{U}$ is a solution and thus $\mathcal{G}^{\prime}[\mathcal{U}]$ should be connected. Hence, $\mathcal{W}$ covers at least $\ell$ edges in $\mathcal{G}$ and thus the Partial Vertex Cover instance is a yes instance.
We also observe that the size of a minimum vertex cover of the subgraph induced by any solution of the reduced Connected Knapsack instance is at most $k+1$ — the set consisting of the vertices in any solution of the Connected Knapsack instance in $\mathcal{G}^{\prime}$ which has weight $1$ and $g$ forms a vertex cover of the subgraph induced by the solution of the Connected Knapsack instance. Also, all the numbers in our reduced Connected Knapsack instance are at most the number of edges of the graph, and Partial Vertex Cover is $\mathsf{W[1]}$-hard parameterized by $k$. Hence, Connected Knapsack is $\mathsf{W[1]}$-hard parameterized by the maximum size of the minimum vertex cover of any solution even when all the numbers are encoded in unary. Therefore, there is no algorithm for Connected Knapsack running in time $\mathcal{O}(f(vcs).\text{poly}(n,s,d))$ unless $\mathsf{ETH}$ fails.
4 Path Knapsack
We now present the results of Path Knapsack. We first show that Path Knapsack is strongly $\mathsf{NP}$-complete by reducing from Hamiltonian Path which is defined as follows.
Definition 8 (Hamiltonian Path)
Given a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ and two vertices $x$ and $y$, compute if there exists a path between $x$ and $y$ which visits every other vertex in $\mathcal{G}$. We denote an arbitrary instance of Hamiltonian Path by $(\mathcal{G},x,y)$.
Hamiltonian Path is known to be $\mathsf{NP}$-complete even for graphs with maximum degree three [11].
Theorem 4.1
Path Knapsack is strongly $\mathsf{NP}$-complete even for graphs with maximum degree three.
Proof
Path Knapsack clearly belongs to $\mathsf{NP}$. To show $\mathsf{NP}$-hardness, we reduce from Hamiltonian Path. Let $(\mathcal{G},x,y)$ be an arbitrary instance of Hamiltonian Path. We consider the following instance $(\mathcal{G},(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d,x,y)$ of Path Knapsack.
$w(u)=0,\alpha(u)=1~{}\forall u\in\mathcal{V},s=0,d=n$
We claim that the two instances are equivalent.
In one direction, let us assume that Hamiltonian Path is yes. Let $\mathcal{P}$ be a Hamiltonian path between $x$ and $y$ in $\mathcal{G}$. Then we have $w(\mathcal{P})=0\leqslant s$ and $\alpha(\mathcal{P})=n\geqslant d$. Hence, the Path Knapsack instance is also a yes instance.
In the other direction, let us assume that the Path Knapsack instance is a yes instance. Let $\mathcal{P}$ be a path between $x$ and $y$ with $w(\mathcal{P})\leqslant s=0$ and $\alpha(\mathcal{P})\geqslant d=n$. Since $\alpha(\mathcal{V})=n=\alpha(\mathcal{P})$ and there is no vertex with zero $\alpha$ value, $\mathcal{P}$ must be a Hamiltonian path between $x$ and $y$. Hence, the Hamiltonian Path instance is also a yes instance.
We observe that all the numbers in our reduced Path Knapsack instance are at most the number of vertices in the graph. Hence, our reduction shows that Path Knapsack is strongly $\mathsf{NP}$-complete.
However, Path Knapsack is clearly polynomial-time solvable for trees, since there exists only one path between every two vertices in any tree.
Observation 1
Path Knapsack is polynomial-time solvable for trees.
One immediate natural question is if Observation 1 can be generalized to graphs of bounded treewidth. The following result refutes the existence of any such algorithm.
Theorem 4.2
[…]
Proof
We show a reduction from Knapsack. Let $(\mathcal{X}=[n],(\theta_{i})_{i\in\mathcal{X}},(p_{i})_{i\in\mathcal{X}},b,q)$ be an arbitrary instance of Knapsack. We consider the following instance $(\mathcal{G}=(\mathcal{V},\mathcal{E}),(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d,x,y)$ of Path Knapsack.
$\mathcal{V}$ =$=\{u_{0}\}\cup\{u_{i},v_{i},w_{i}:i\in[n]\}$
$\mathcal{E}$ =$=\{\{u_{i},v_{i+1}\},\{u_{i},w_{i+1}\}:0\leqslant i\leqslant n-1\}\cup\{\{u_{i},v_{i}\},\{u_{i},w_{i}\}:i\in[n]\}$
$w(v_{i})$ $=\theta_{i},\alpha(v_{i})=p_{i}~{}\forall i\in[n],w(u_{i})=\alpha(u_{i})=0~{}\forall i\in[n]_{0},w(w_{i})=\alpha(w_{i})=0~{}\forall i\in[n]$
$s$ $=b,d=q,x=u_{0},y=u_{n}$
Figure 1 shows a schematic diagram of the reduced instance. We now claim that the two instances are equivalent.
Figure 1: Reduced instance of Path Knapsack in Theorem 4.2.
In one direction, let us suppose that the Knapsack instance is a yes instance. Let $\mathcal{I}\subseteq\mathcal{X}$ be such that $\sum_{i\in\mathcal{I}}\theta_{i}\leqslant b$ and $\sum_{i\in\mathcal{I}}p_{i}\geqslant q$. We now consider the path $\mathcal{P}$ from $u_{0}$ to $u_{n}$ consisting of $2n$ edges where the $2i$-th vertex (from $u_{0}$) is $v_{i}$ if $i\in\mathcal{I}$ and $w_{i}$ otherwise for every $i\in[n]$; the $(2i-1)$-th vertex (from $u_{0}$) is $u_{i-1}$ for every $i\in[n]$ and the last vertex is $u_{n}$. Clearly, we have $w(\mathcal{P})=\sum_{i\in\mathcal{I}}\theta_{i}\leqslant b=s$ and $\alpha(\mathcal{P})=\sum_{i\in\mathcal{I}}p_{i}\geqslant q=d$. Hence, the Path Knapsack instance is a yes instance.
In the other direction, let us assume that the Path Knapsack instance is a yes instance. Let $\mathcal{P}$ be a path between $x$ and $y$ with $w(\mathcal{P})\leqslant s=b$ and $\alpha(\mathcal{P})\geqslant d=q$. We observe that all paths from $u_{0}$ to $u_{n}$ in $\mathcal{G}$ have $2n$ edges; $(2i-1)$-th vertex is $u_{i-1}$ and $2i$-th vertex is either $v_{i}$ or $w_{i}$ for every $i\in[n]$ and the last vertex is $u_{n}$. We consider the set =$\mathcal{I}=\{i\in[n]:\text{the $(2i-1)$-th vertex in $\mathcal{P}$ is $v_{i}$}\}\subseteq\mathcal{X}$. We now have $\sum_{i\in\mathcal{I}}\theta_{i}=w(\mathcal{P})\leqslant s=b$ and $\sum_{i\in\mathcal{I}}p_{i}=\alpha(\mathcal{P})\geqslant d=q$. Hence, the Knapsack instance is a yes instance.
We also observe that the pathwidth of $\mathcal{G}$ is at most two (there is a path decomposition with bag size being at most three) which proves the result.
Theorem 4.2 leaves the following question open: does there exist an algorithm for Path Knapsack which runs in time $\mathcal{O}(f(\text{tw})\cdot\text{poly}(n,s,d))$? We answer this question affirmatively in the following result. The dynamic programming algorithm is very similar to Theorem 3.3.
Theorem 4.3
[…]
Using the technique in Theorem 3.4, we use Theorem 4.3 as a black-box to obtain the following approximation algorithm. We again omit its proof due to its similarity with Theorem 3.4.
Theorem 4.4
[…]
We next consider the size of the minimum vertex cover $vcs$ of the subgraph induced by a solution $\mathcal{W}\subseteq\mathcal{V}[\mathcal{G}]$. We observe that the size of the minimum vertex cover of $\mathcal{G}[\mathcal{W}]$ is at least half of $|\mathcal{W}|$ since there exists a Hamiltonian path in $\mathcal{G}[\mathcal{W}]$. Hence, it is enough to design an $\mathsf{FPT}$ algorithm with parameter $|\mathcal{W}|$. Our algorithm is based on the color coding technique [9].
Theorem 4.5
There is an algorithm for Path Knapsack running in time $\mathcal{O}\left((2e)^{k}k^{\mathcal{O}(\log k)}n^{\mathcal{O}(1)}\right)$ where $k$ is the number of vertices in the solution.
Proof
Let $(\mathcal{G}=(\mathcal{V},\mathcal{E}),(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d,x^{\prime},y^{\prime})$ be an arbitrary instance of Path Knapsack and $k$ the number of vertices in the solution. We can assume without loss of generality that we know $k$ since there are only $n-1$ possible values of $k$ namely $2,3,\ldots,n$ — we can simply run the algorithm starting from $k=1$, incrementing $k$ in every iteration, till we obtain a solution which we output. We color every vertex uniformly randomly from a palette of $k$ colors independent of everything else. Let $\chi:\mathcal{V}\longrightarrow[k]$ be the resulting coloring. For every non-empty subset $S\subseteq[k],S\neq\emptyset$ and vertex $x\in\mathcal{V}$, we define a boolean variable PATH$(S,x)$ to be true if there is a path which starts from $x^{\prime}$, ends at $x$, and contains exactly one vertex of every color in $S$; we call such a path $S$-colorful. If PATH$(S,x)$ is true, then we also define $D[S,x]=\{(w,\alpha):\exists\text{ an $x^{\prime}$ to $x$ $S$-colorful path $\mathcal{P}$ such that }w(\mathcal{P})=w,\alpha(\mathcal{P})=\alpha\text{ for every other $x^{\prime}$ to $x$ $S$-colorful path $\mathcal{Q}$, we have either }w(\mathcal{Q})>w\text{ or }\alpha(\mathcal{Q})<\alpha\}$. For $|S|=1$, we note that PATH$(S,x^{\prime})$=true if and only if $\{\chi(x^{\prime})\}=S$; PATH$(S,x)$=false for every $x\in\mathcal{V}\setminus\{x^{\prime}\}$ and $|S|=1$; $D[S,x^{\prime}]=\{(w(x^{\prime}),\alpha(x^{\prime}))\}$ if $\{\chi(x^{\prime})\}=S$; $D[S,x]=\emptyset$ for every $x\in\mathcal{V}\setminus\{x^{\prime}\}$ and $S\subseteq\mathcal{V}$. We update $\text{PATH}(S,x)$ as per the following recurrence for $|S|>1$.
$\text{PATH}(S,x)=\begin{cases}\bigvee\{\text{PATH}(S\setminus\{\chi(x)\},v):\{v,x\}\in\mathcal{E}\}&\text{if }\chi(x)\in S\\ \text{{\sc false}}&\text{otherwise}\end{cases}$
When we update any $\text{PATH}(S,x)$ to be true, we update $D[S,x]$ as follows. For every $\{x,y\}\in\mathcal{E}$ if $\text{PATH}(S\setminus\{\chi(x)\},y)$ is true, then we do the following: for every $(w,\alpha)\in D[S\setminus\{\chi(x)\},y]$, we put $(w+w(x),\alpha+\alpha(x))$ in $D[S,x]$ if $w+w(x)\leqslant s$; and we finally remove all dominated pairs from $D[S,x]$. We output yes if $\text{PATH}([k],y^{\prime})$ is true and there exists a $(w,\alpha)\in D[[k],y^{\prime}]$ such that $w\leqslant s$ and $\alpha\geqslant d$. Otherwise, we output no.
Proof of correctness: If there does not exist any colorful path between $x$ and $y$ of length $k$, then the algorithm clearly outputs no. Suppose now that the instance is a yes instance. Then there exists a colorful path $\mathcal{P}=(u_{1}(=x),u_{2},\ldots,u_{k}(=y))$ between $x$ and $y$ such that $w(\mathcal{P})\leqslant s$ and $\alpha(\mathcal{P})\geqslant d$. Let us define $S_{i}=\{u_{j}:j\in[i]\}$ for $i\in[k]$. Then $\text{PATH}(S_{i},u_{i})$ is true and $D[S_{i},u_{i}]$ either contains $(w(S_{i}),\alpha(S_{i}))$ or any pair which dominates $(w(S_{i}),\alpha(S_{i}))$ for every $i\in[k]$. Hence, the algorithm outputs yes.
Running time analysis: If there exists a colorful path between $x$ and $y$, then the algorithm finds it in time $\mathcal{O}\left(2^{k}n^{\mathcal{O}(1)}\right)$. If there exists a path between $x$ and $y$ containing $k-1$ edges (that is, $k$ vertices including $x$ and $y$), then one such path becomes colorful in the random coloring with probability at least
$\frac{k!}{k^{k}}\geqslant e^{-k}.$
Hence, by repeating $\mathcal{O}(e^{k})$ times and outputting yes if any run of the algorithm outputs yes, the above algorithm achieves a success probability of at least $2/3$. We can use the splitters to derandomize the algorithm above to obtain a deterministic algorithm for Path Knapsack which runs in time $\mathcal{O}\left((2e)^{k}k^{\mathcal{O}(\log k)}n^{\mathcal{O}(1)}\right)$ [9, Section 5.6.2].
Theorem 4.5 immediately implies the following result.
Corollary 2
There is an algorithm for Path Knapsack running in time $\mathcal{O}\left((2e)^{2vcs}vcs^{\mathcal{O}(\log vcs)}n^{\mathcal{O}(1)}\right)$ where $vcs$ is the size of the minimum vertex cover of the subgraph induced by the solution.
5 Shortest Path Knapsack
We now consider Shortest Path Knapsack. We observe that all the $u$ to $v$ paths in the reduced instance of Path Knapsack in the proof of Theorem 4.2 are of the same length. Hence, we immediately obtain the following result as a corollary of Theorem 4.2.
Corollary 3
[…]
Interestingly, Shortest Path Knapsack admits a pseudo-polynomial-time algorithm for any graph unlike Connected Knapsack and Path Knapsack.
Theorem 5.1
There is an algorithm for Shortest Path Knapsack running in time $\mathcal{O}(n\log n\cdot\min\{s^{2},(\alpha(\mathcal{V}))^{2}\})$, where $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is the input graph.
Proof
Let $(\mathcal{G}=(\mathcal{V},\mathcal{E}),(c(e))_{e\in\mathcal{E}},(w(u))_{u\in\mathcal{V}},(\alpha(u))_{u\in\mathcal{V}},s,d,x,y)$ be an arbitrary instance of Shortest Path Knapsack. We design a greedy and dynamic-programming based algorithm. For every vertex $v\in\mathcal{V}$, we store a boolean marker $b_{v}$, the distance $\delta_{v}$ of $v$ from $x$ and a set $D_{v}=\{(w,\alpha):\exists$ an $x$ to $v$ shortest-path $\mathcal{P}$ such that $w(\mathcal{P})=w,\alpha(\mathcal{P})=\alpha$ and for every other $x$ to $v$ shortest-path $\mathcal{Q}$, we have either $w(\mathcal{Q})>w$ or $\alpha(\mathcal{Q})<\alpha\}$. That is, we store undominated weight-value pairs of all shortest $x$ to $v$ paths. We initialize $b_{x}=\text{{\sc false}},\delta_{x}=0,D_{x}=\{(w(x),\alpha(x)\},b_{u}=\text{{\sc false}},\delta_{u}=\infty,$ and $D_{u}=\emptyset$ for every $u\in\mathcal{V}\setminus\{x\}$.
Updating DP table: We pick a vertex $z=\operatorname*{arg\,min}_{v\in\mathcal{V}:b_{v}=\text{{\sc false}}}\delta_{v}$. We set $b_{z}=\text{{\sc true}}$. For every neighbor $u$ of $z$, we set $\delta_{u}=\min\{\delta_{u},\delta_{z}+c(\{z,u\})\}$. If $\delta_{u}=\delta_{z}+c(\{z,u\})$, then update $D_{u}$ as follows: for every $(w,\alpha)\in D_{z}$, we put $(w+w(u),\alpha+\alpha(u))$ in $D_{u}$ if $w+w(u)\leqslant s$; and we finally remove all dominated pairs from $D_{u}$ just before finishing each iteration. If we have $b_{v}=\text{{\sc true}}$ for every vertex, then we output yes if there exists a pair $(w,\alpha)\in D_{y}$ such that $w\leqslant s$ and $\alpha\geqslant d$. Else, we output no.
We now argue the correctness of our algorithm. Following the proof of correctness of the classical Dijkstra’s shortest path algorithm, we observe that if $b_{v}$ is true for any vertex $v\in\mathcal{V}$, its distance from $x$ is $\delta_{v}$ [8]. We claim that at the end of updating a table entry in every iteration, the following invariant holds: for every vertex $v\in\mathcal{V}$ such that $b_{v}=\text{{\sc true}}$ and every $x$ to $v$ undominated shortest path $\mathcal{P}$ using only vertices marked true, we have $(w(\mathcal{P}),\alpha(\mathcal{P}))\in D_{v}$. The invariant holds vacuously after initialization since every vertex is marked false. Let us assume that the invariant hold after $i$ iterations; $\mathcal{V}_{T}$ be the set of vertices which are marked true after $i$ iterations. Suppose the algorithm picks the vertex $z_{i+1}$ in the $(i+1)$-th iteration. Let $\mathcal{P}^{*}=x,\ldots,y,z_{i+1}$ be an undominated $x$ to $z_{i+1}$ shortest path using the vertices marked true only. Then we claim that the prefix of the path $\mathcal{P}^{*}$ from $x$ to $y$, let us call it $\mathcal{Q}=x,\ldots,y$, is an undominated $x$ to $y$ shortest path. That $\mathcal{Q}$ is a shortest path from $x$ to $y$ follows from standard proof of correctness of Dijkstra’s algorithm [8]. Now, to show that $\mathcal{Q}$ is undominated, let us assume that another shortest path $\mathcal{R}$ from $x$ to $y$ dominates $\mathcal{Q}$. Then the shortest path $\mathcal{R}^{\prime}$ from $x$ to $z_{i+1}$ which is $\mathcal{R}$ followed by $z_{i+1}$ also dominates $\mathcal{P}^{*}$ contradicting our assumption that $\mathcal{P}^{*}$ is an undominated shortest path from $x$ to $z_{i+1}$. Hence, the pair $(w(\mathcal{P}^{*}),\alpha(\mathcal{P}^{*}))$ is put in $D_{z_{i+1}}$ in the $(i+1)$-th iteration. And since the algorithm only removes dominated pairs in every iteration, the pair $(w(\mathcal{P}^{*}),\alpha(\mathcal{P}^{*}))$ remains in $D_{z_{i+1}}$ after the end of $(i+1)$-th iteration also. This proves that the invariant holds after $(i+1)$-th iteration also, and thus the correctness of the algorithm. For every vertex $v\in\mathcal{V}$, the cardinality of $D_{v}$ is at most $s$ and also at most $\alpha(\mathcal{V})$ and thus at most $\min\{s,\alpha(\mathcal{V})\}$. Implementing the above algorithm using a standard Fibonacci heap-based priority queue to find $\operatorname*{arg\,min}_{v\in\mathcal{V}:b_{v}=\text{{\sc false}}}\delta_{v}$ gives us a running time $\mathcal{O}(n\log n\cdot\min\{s^{2},(\alpha(\mathcal{V}))^{2}\})$.
Clearly, Shortest Path Knapsack admits a polynomial-time algorithm for trees since only one path exists between every two vertices.
Observation 2
Shortest Path Knapsack is in $\mathsf{P}$ for trees.
Using the technique in Theorem 3.4, we use Theorem 5.1 as a black-box to obtain the following approximation algorithm. We again omit its proof due to its similarity with Theorem 3.4.
Theorem 5.2
[…]
6 Conclusion
We study the classical Knapsack problem with various graph theoretic constraints, namely connectedness, path and shortest path. We show that Connected Knapsack and Path Knapsack are strongly $\mathsf{NP}$-complete whereas Shortest Path Knapsack admits a pseudo-polynomial time algorithm. All the three problems admit FPTASes for bounded treewidth graphs; only Shortest Path Knapsack admits an FPTAS for arbitrary graphs. It would be interesting to explore if meta-theorems can be proven in this general theme of knapsack on graphs.